OpenStudy (anonymous):

A circle is drawn through the vertices of a square with sides of 5 inches. If a dart is thrown at random into the circle, what is the probability it will NOT lie in the square. Use pythagoreean theorem to find the radius of the circle.

OpenStudy (ybarrap):

R= Radius = (1/2) of diagonal of square = (1/2)[ 5^2 + 5^2 ]] ^ (1/2) = sqrt(50)/2 P (not in square) = (Area circle - Area square) / Area circle = (pi*R^2 - 25)/(pi*R^2) = (pi*(sqrt(50)/2)^2 - 25)/(pi*(sqrt(50)/2)^2) = (pi*(50/4) - 25)/(pi*50/4) =25*(2*pi/4 - 1)/((2*pi/4)*25) = (pi/2 - 1)/(pi/2) = 1 - 1/(pi/2) = 1 - 2/pi